A study of the effects of exercise used rats bred to have high or low capacity for exercise. There were 13 high-capacity and 24 low-capacity rats. To compare the mean blood pressure of the two types of rats using the conservative Option 2 t-procedures. What are the associated degrees of freedom?
Answer #1

The degree of freedom are

n1 - 1 = 13 - 1 = 12

n2 - 1 = 24 - 1 = 23